there are 2 linear equations, one is y=6-3/4x the other is y=x-3 at what point do these lines intersect
Answer by funmath(2933)

there are 2 linear equations, one is y=6-3/4x the other is y=x-3 at what point do these lines intersect
Let their y's =.
6-3/4x=x-3
4(6-3/4x)=4(x-3)
24-12/4x=4x-12
24-3x=4x-12
24-3x+3x=4x+3x-12
24=7x-12
24+12=7x-12+12
36=7x
36/7=7x/7
x=36/7
Substitute this in for x in one of the equations and solve for y:
y=(36/7)-3
y=36/7-21/7
y=15/7
The point of intersection is (x,y)=(36/7,15/7)
Happy Calculating!!!